y=x+1
y=-2x-4 please help??????

Found 2 solutions by vidhyak, checkley71:
Answer by vidhyak(98)

y=x+1 --------equation 1
y=-2x-4 --------equation 2
Subtract eq 1 and eq 2
Y term gets cancelled and you get
(y-y) = x -(-2x) + 1-(-4)
0 = x+2x +1+4
3x + 5 = 0
3x = -5
x = -5/3
Substitute the value of x = -5/3 in eq 1.
y = x + 1
y = (-5/3) + 1
y = -2/3





Answer by checkley71(8403)

IF Y=X+1 & Y=-2X-4 THE X+1=-2X-4 OR 3X=-5 OR X=-5/3 OR Y=-5/3+1 Y=-5/3+3/3 OR Y=-2/3
PROOF -2/3=-2*(-5/3)-4 OR -2/3=10/3-4 OR -2/3=10/3-12/3 OR -2/3=-2/3